DIRC Operations Weekly Report
Period: Wednesday 2007/3/27 - Wednesday 2007/4/4
Oncall: Bradley Wogsland
Fast Monitoring: Alejandro Perez, Bradley Wogsland
PEP peak lumi: ~8.3 x 1033
(previous week: ~9 x 1033)
BaBar integrated lumi:1.97 fb-1
(previous week: 3.41 fb-1)
Run 6 PEP delivery:2.07 fb-1
(previous week: 3.43 fb-1)
Run 6 total delivery: ~22 fb-1
Run 6 total BaBar recorded: 21.2 fb-1
including 0 fb-1 off-peak data
PEP activities this week:
LCLS activities this week:
Broken water pipe shut down BaBar's cooling water 4/3
Foreseen work opportunities (PEP down periods):
8 hr MD planned for Wednesday
BaBar activities/issues this week:
Came back up fairly well from the ROD
- Nicolas added a BCS SIAM alarm in the ALH after noticing we forgot to
reinhibit it 3/29.
- Updated DQM short term instructions 3/30 because the flaky board in
sector 8 was showing up on the JAS plots.
- Paged 3/31 for high background (scalars all in red) by pilot who
claimed MCC was unresponsive to him. Called MCC and they got the scalars
back in the green in 15 mins. Told pilot to be more persistant in the
future - which seemed to work for the rest of the weekend.
logbook for details
- Replace two DFBs in sector 1 with a group of 8 innefficient or
flaky channels (require bridge).
- Replace new DFB in sector 8 with 8 often dead channels and figure
out what is killing DFBs in that slot.
Jose, Alejandro and Bradley rotating backup & on-call commissioning